8,9and 10 also please draw cash flow diagrams
customer service department. I he company can earn an interest at 10% on the lump sum deposited now and it wishes to withdraw the money in the following increments. .Year 1: $25,000 to purchase a computer Year 2: $3000 to purchase additional hardware . Year 3: No expenses How much money must be deposited now to cover anticipated expenses over next 4 years? ANNUAL WORTH METHOD Year 4: $5000 to purchase s/w upgrades. 7. -Suppose you make an annual contribution of $3000 to your savings accounit at the end of each year for 10 years. If the account earns 7% interest annually, how much can be withdrawn at the end of 10 years? Suppose the ten deposits were made at the beginning of each perlod, what is the balance at the end of period 10? 8. Consider 3 ¡vestment plans for an individual who just celebrated his 24th birthday. (1-8%) Invest $2000 per year for first 10 years. At the end of 10 years make no further Investments, but reinvest the amount for next 30 years. Do nothing for first 10 years. Then start investing $2000/year for the next 30 years. L. Il. Ill. Invest $2000/ year for the entire 40 years Note that all investments are made at the birthday of each year, the first deposit will be made on 25 birthday (n-1), Calculate the balance as on 64 birthday (n-40) Blogen Co., has borrowed S250,000 to purchase lab equipment for gene splicing at an interest rate of 8% and is to be repald In equal Instalments over next 9. 6years. I. Compute the annual Instalment ll. Suppose Blo Gen wants to negotiate with the bank to defer the first loan payment until the end of year 2 (but stil desires to pay the 6 Instalments at 8% interest) what should be the annual Instalment ( deferred annuity) 10. You plan to retire 33 years from now. You expect that you will live 27 years after retiring. You want to have enough money upon reaching retirement age to withdraw Rs. 1,80,000 from the account at the beginning of each year you expect to live, and yet still have Rs. 25,00,000 left in the account at the time of your expected death ( 60 years from now). You plan to accumulate the retirement fund by making equal annual deposits at the end of each year for the nest 33 years. You expect that you will be able to earn 12% per year on your own deposits. However, you only expect to earn 6% per year on your investment after you retire since you wiG
choose to place the money in less risky investments. What equal annual deposits must you make each year to reach your retirement goal? GRADIENT SERIES: 11. A textile mill has just purchased a lift truck that has a useful life of 5 years. The engineer estimates that the maintenance costs for the truck during the first year will be $1000. As the truck ages, the maintenance costs are expected to increase at a rate of $250 per year over the remaining life. The firm wants to set up a maintenance account that earns 12% annual interest. All future maintenance expenses will be paid out of this account. How much does the firm have to deposit in the account now? 12. Suppose that you make a series of annual deposits into a bank account that pays 10% interest. The initial deposit at the end of 1t year is $1,200. The deposit amounts decline by $200 in each of the next 4 years. How much would you have immediately after the fifth deposit? 13. John and Barbara have just opened 2 savings accounts at their bank. John wants to deposit $1000 in his account at the end of the first year and Increase the amount by $300 for each of the next 5 years. Barbara wants to deposit an equal amount each year for next 6 years. What should be Barbaras annual deposit so that the twaccounts
Answer #1

8)

I) Future Value of Annuity due

F = A[(1+r)^n - 1]/r(1+r)

Here A = 2000

r = 8%

n = 10

F = 2000*[(1.08)^10 - 1]/0.08*1.08 = 31290.97

This future value will be reinvested for next 30 years

Final Future Value = 31290.97(1.08)^30 = 314870.30

ii)

F = 2000*[(1.08)^30 - 1]/0.08*1.08 = 244691.7

iii)

F = 2000*[(1.08)^40 - 1]/0.08*1.08 = 559562.1